The time, t, it takes for an ice cube to melt is inversely proportional to the temperature of the water it is in. If it takes an ice cube 2.5 minutes to melt in 70 degree F water, how long will it take an ice cube of the same size to melt in 50 degree water?
Found 2 solutions by richwmiller, Alan3354:
Answer by richwmiller(17219)

y=k/x
2.5=k/70
2.5*70=k
175=k
y=175/50=7/2=3.5 minutes

Answer by Alan3354(69443)

The best answer is insufficient data.
70 F is not 1.4 times 50 F
Any calculations that multiply temperature have to use "absolute zero" as reference.
--------------
70F = 21.11C
50F = 10C
That makes the ratio 2.11 instead of 1.4. Now what?
----------------
The author of the problem doesn't know what he/she is talking about.